# Gorenstein cohomological dimension and stable categories for groups

### v4

@@ 1. Introduction
The origin of Gorenstein projective dimension may date back to the study of $G$-dimension by Auslander and Bridger [2] in 1960s. Let $G$ be any group. Over the group ring $RG$, the coefficient ring $R$ is an $RG$-module with trivial group action. The cohomological dimension ${\rm cd}_{R}G$, and the Gorenstein cohomological dimension ${\rm Gcd}_{R}G$, is defined as the projective dimension and Gorenstein projective dimension of the trivial $RG$-module $R$ respectively. Studying groups through their cohomological dimensions arose from both topological and algebraic sources; see [9]. We refer to Introduction of [22] for more details on the geometric significance of the Gorenstein cohomological dimension of groups.
@@
The “Gcd” can be considered as an assignment of invariants for the pairs of groups and coefficient rings $(G,R)$. In Section 3 and 4, we study the properties of Gcd under changes of groups and coefficient rings, respectively. we define an order $(H,S)\leq(G,R)$ for such pairs; see Definition 3.1. Inspired by [22, 40], we use Theorem 2.4 to show that when $R$ is a commutative Gorenstein regular ring, $\mathrm{Gcd}_{R}H\leq\mathrm{Gcd}_{R}G$ if $(H,R)\leq(G,R)$, and $\mathrm{Gcd}_{S}G\leq\mathrm{Gcd}_{R}G$ if $(G,S)\leq(G,R)$; see Proposition 3.2 and Proposition 4.1. These generalize [22, Propositions 2.1 and 2.4] and [40, Theorem 3.2]. Consequently, “Gcd” preserves the order of pairs of groups and commutative Gorenstein regular rings; see Corollary 4.3.
By Serre’s Theorem, ${\rm cd}_{\mathbb{Z}}G={\rm cd}_{\mathbb{Z}}H$ for a torsion-free group $G$ and any subgroup $H$ of finite index; see e.g. [9, Theorem VIII 3.1]. We have a Gorenstein version of Serre’s Theorem for any group and its subgroups of finite index over any coefficient ring; see Theorem 3.3. Note that the equality ${\rm Gcd}_{R}G={\rm Gcd}_{R}H$ was stated in [22, Corollary 2.10] under assumptions that the coefficient ring $R$ is of finite weak global dimension and $H$ is a normal submodule of $G$.
@@
For model category $\mathcal{F}ib$, the associated homotopy category $\mathrm{Ho}(\mathcal{F}ib)$ is obtained by formally inverting the weak equivalences. It is standard that $\mathrm{Ho}(\mathcal{F}ib)$ is triangle equivalent to the stable category $\underline{\mathcal{C}of}(RG)$; see Corollary 6.4. Moreover, ${\rm Ho}(\mathcal{F}ib)$ is equivalent to the stable module category ${\rm StMod}(RG)$; see Theorem 6.8. It is worth to remark that the modules of ${\rm StMod}(RG)$ was assumed to be countably presented in [6], and Benson suggested that one can lift the countability hypothesis based on [6, Open Problem 2]. In fact, the objects of ${\rm StMod}(RG)$ are precisely the fibrant $RG$-modules, and the countability assumption of modules in ${\rm StMod}(RG)$ is not necessary.
In [34, Theorem 3.10], some equivalences of triangulated categories
@@ 2. Finiteness of Gorenstein cohomological dimension of groups
Let $M$ be a Gorenstein projective $RG$-module. Then there is a totally acyclic complex of projective $RG$-modules $\cdots\rightarrow P_{1}\rightarrow P_{0}\rightarrow P_{-1}\rightarrow\cdots$ such that $M\cong\mathrm{Ker}(P_{0}\rightarrow P_{-1})$. Since any projective $RG$-module is also $R$-projective, by restricting this totally acyclic complex, we get an acyclic complex of projective $R$-module. Noting that $R$ is left Gorenstein regular, every projective $R$-module $P$ has finite injective dimension. Apply ${\rm Hom}_{R}(-,P)$ to the above complex, and it is easy to see that the acyclic complex of projective $R$-module is totally acyclic by induction on the injective dimension of $P$. Hence, $M$ is also a Gorenstein projective $R$-module. ∎
@@
Let ${\rm Gcd}_{R}G={\rm Gpd}_{RG}R=n$. It follows from [30, Theorem 2.10] that there exists an exact sequence $0\rightarrow K\rightarrow M\rightarrow R\rightarrow 0$, where $M$ is a Gorenstein projective $RG$-module, and ${\rm pd}_{RG}K=n-1$. For $M$, there is an exact sequence of $RG$-modules $0\rightarrow M\rightarrow P\rightarrow L\rightarrow 0$, where $L$ is Gorenstein projective and $P$ is projective. We consider the following pushout of $M\rightarrow R$ and $M\rightarrow P$:
@@
From the middle row we infer that ${\rm pd}_{RG}\Lambda={\rm pd}_{RG}K+1=n$. It follows from Lemma 2.1 that $L$ is also a Gorenstein projective $R$-module, and then the sequence $0\rightarrow R\rightarrow\Lambda\rightarrow L\rightarrow 0$ is $R$-split. Moreover, as an $R$-module, $\Lambda\cong L\oplus R$ is Gorenstein projective. By [27, Proposition 10.2.3], which says that projective dimension of any Gorenstein projective module is either zero or infinity, we induce from ${\rm pd}_{R}\Lambda\leq{\rm pd}_{RG}\Lambda=n$ that $\Lambda$ is a projective $R$-module. This completes the proof. ∎
@@
It is well known that for any finite group $G$ and any coefficient ring $R$, ${\rm Gcd}_{R}G=0$ is finite; see [1, Proposition 2.19] for $R=\mathbb{Z}$, [22, Corollary 2.3] for any commutative ring $R$, and Proposition 3.4 below for any associative ring $R$.
@@ 3. Gorenstein cohomological dimension under changes of groups
By Serre’s theorem, if $G$ is a torsion-free group and $H$ is a subgroup of finite index, then $\mathrm{cd}_{\mathbb{Z}}H=\mathrm{cd}_{\mathbb{Z}}G$; see e.g. [38, Theorem 9.2] or [9, Theorem VIII 3.1]. The following result might be regarded as a Gorenstein version of Serre’s theorem. We remark that the equality for the coefficient ring $\mathbb{Z}$ holds by [1, Proposition 2.18], and the result was also proved in [22, Corollary 2.10] under the assumptions that the coefficient ring $R$ is of finite weak global dimension and $H$ is a normal submodule of $G$.
@@
that $\mathrm{Ext}_{RH}^{1}(L_{0},Q)=0$. Consider $0\rightarrow K\stackrel{{\scriptstyle d}}{{\rightarrow}}P_{0}\rightarrow L_{0}\rightarrow 0$ as an exact sequence of $RH$-modules, where $K$ is Gorenstein projective and $P_{0}$ is projective. By [30, Corollary 2.11], $L_{0}$ is a Gorenstein projective $RH$-module. Hence, for any projective $RG$-module $P$, we infer from ${\rm Ext}^{1}_{RG}(L_{0},{\rm Ind}_{H}^{G}P)\cong{\rm Ext}_{RH}^{1}(L_{0},P)=0$ that ${\rm Ext}^{1}_{RG}(L_{0},P)=0$. This implies that the sequence $0\rightarrow K\rightarrow P_{0}\rightarrow L_{0}\rightarrow 0$ remains exact after applying ${\rm Hom}_{RG}(-,P)$.
Then, repeat the above argument for $L_{0}$, we will obtain inductively an acyclic complex $0\rightarrow K\rightarrow P_{0}\rightarrow P_{-1}\rightarrow P_{-2}\rightarrow\cdots$ with each $P_{i}$ a projective $RG$-module, which remains acyclic after applying ${\rm Hom}_{RG}(-,P)$ for any projective $RG$-module $P$. Recall that ${\rm Ext}^{i}_{RG}(K,P)=0$. Consequently, we infer from [30, Proposition 2.3] that $K$ is a Gorenstein projective $RG$-module, as desired. ∎
@@
(2$\Longrightarrow$(3) is trivial.
(3$\Longrightarrow$(1) Consider the subgroup $H=\{1\}$ of $G$. Conversely, let us assume that the group $G$ is not finite. We claim that for any $\mathbb{Z}$-module $M$, the $G$-invariant submodule $(\mathrm{Ind}_{H}^{G}M)^{G}$ of the $\mathbb{Z}G$-module $\mathrm{Ind}_{H}^{G}M=\mathbb{Z}G\otimes_{\mathbb{Z}}M$ equals to 0. By [9, Proposition III 5.1], $\mathrm{Ind}_{H}^{G}M$ is the direct sum of the transform $gM$, where $g$ is an element of $G$ other than the identity. Consider an arbitrary element $x=\sum g_{i}m_{i}$ with all but finitely many $m_{i}\neq 0$. Without loss of generality, let $m_{1}\neq 0$. We may take the summand entry $g_{1}m_{1}$ of $x$ and a summand $g^{\prime}M$ with $g^{\prime}\neq g_{i}$ for those $g_{i}m_{i}\neq 0$. Then, there exists $g^{\prime\prime}\in G$ such that $g^{\prime\prime}(g_{1}m_{1})=g^{\prime}m_{1}$ because the $G$-action is transitive on the summand. Thus $x$ is not fixed by $G$, and so $(\mathrm{Ind}_{H}^{G}M)^{G}=0$, as claimed.
Since ${\rm Gcd}_{\mathbb{Z}}G=0$, it follows from Lemma 2.2 that there exists a $\mathbb{Z}$-split $\mathbb{Z}G$-monomorphism $\mathbb{Z}\rightarrow\Lambda$, where $\Lambda$ is $\mathbb{Z}$-free and $\mathbb{Z}G$-projective. Then, we get $(\mathrm{Ind}_{H}^{G}\Lambda)^{G}=\mathrm{Hom}_{\mathbb{Z}G}(\mathbb{Z},\mathbb{Z}G\otimes_{\mathbb{Z}}\Lambda)\neq 0$, which conflicts with the above assertion $(\mathrm{Ind}_{H}^{G}M)^{G}=0$. Hence, the contradiction forces $G$ to be a finite group.
@@
Let $G$ be a group, $H$ be a finite subgroup of $G$, and denote by $N_{G}(H)$ the normalizer of $H$ in $G$. The Weyl group of $H$ is $W=N_{G}(H)/H$. Note that modules over the Weyl group $W=N_{G}(H)/H$ are exactly the ones over the group $G$ on which the subgroup $H$ acts trivially. That is, for any commutative coefficient ring $R$, the class of $RW$-modules contains precisely all those $RG$-modules $M$ with $M^{H}=M$.
@@
Let $R$ be a commutative Gorenstein regular ring, $G$ be a group. For any finite subgroup $H$, we have ${\rm Gcd}_{R}(N_{G}(H)/H)\leq{\rm Gcd}_{R}G$.
@@ 4. Extension of coefficient rings
Since the inequality obviously holds if $\mathrm{Gcd}_{R}G=\infty$, it only suffices to consider the case where $\mathrm{Gcd}_{R}G=n$ is finite. In this case, it follows immediately from Lemma 2.2 that there exists an $R$-split monomorphism of $RG$-modules $\iota:R\rightarrow\Lambda$, where $\Lambda$ is $R$-projective and $\mathrm{pd}_{RG}\Lambda=n$. Note that $S$ is an extension ring of $R$, then $S$ is a both a trivial $SG$-module and an $R$-module. By applying $S\otimes_{R}-$, we have an $S$-split monomorphism of $SG$-modules $S\otimes_{R}\iota:S\rightarrow S\otimes_{R}\Lambda$, where $S\otimes_{R}\Lambda$ is $S$-projective since $\Lambda$ is $R$-projective.
@@
Since $S\otimes_{R}P_{i}$ are projective $SG$-modules, we have $\mathrm{pd}_{SG}(S\otimes_{R}\Lambda)\leq n$. For the trivial $SG$-module $S$, it follows immediately from [22, Proposition 1.4] that $\mathrm{Gpd}_{SG}S\leq\mathrm{pd}_{SG}(S\otimes_{R}\Lambda)$. Then, $\mathrm{Gcd}_{S}G\leq\mathrm{Gcd}_{R}G$ holds as expected. ∎
Specifically, considering the pairs $(G,R)\leq(G,\mathbb{Z})$ and $(G,\mathbb{Q})\leq(G,\mathbb{Z})$, we reobtain [22, Propostion 2.1] and [40, Theorem 3.2] respectively.
@@
If $(H,S)\leq(G,R)$, then we have both $(H,S)\leq(G,S)$ and $(G,S)\leq(G,R)$. By Proposition 3.2, it follows that ${\rm Gcd}_{S}H\leq{\rm Gcd}_{S}G$. By Proposition 4.1, $\mathrm{Gcd}_{S}G\leq\mathrm{Gcd}_{R}G$ holds. Then, we deduce the desired inequality $\mathrm{Gcd}_{S}H\leq\mathrm{Gcd}_{R}G$. ∎
@@
Note that the second inequality follows from Corollary 4.2. The first inequality obviously holds if ${\rm Gcd}_{R}G$ is infinite. Furthermore, if we assume the finiteness of ${\rm Gcd}_{R}G$, then the desired inequality follows immediately from Lemmas 2.2 and 2.3. ∎
@@ 5. Cofibrant modules and Gorenstein projective modules
Recall that a group $G$ is of type $\Phi_{R}$, if it has the property that an $RG$-module is of finite projective dimension if and only if its restriction to any finite subgroup is of finite projective dimension; see [34, Definition 2.1]. It was conjectured that a group $G$ is of type $\Phi_{\mathbb{Z}}$ if and only if $G$ admits a finite dimensional model for $\underline{E}G$, the classifying space for the family of the finite subgroups of $G$; the sufficient condition was proved to be true. We refer to [39] for the details.
@@
Let $G$ be an $\mathrm{LH}\mathfrak{F}$-group, $R$ a commutative ring. For any Gorenstein projective $RG$-module $M$, if ${\rm pd}_{R}M<\infty$, then $M$ is cofibrant.
@@
We might use Propositions 5.6 and 5.7 to get the following, which supports [19, Conjecture A] and recovers [19, Corollary C], [7, Theorem 3.11] and [8, Theorem 3.4].
@@
If $G$ is of type $\Phi_{R}$, then by [6, Lemma 3.4(ii)] we have $\mathrm{pd}_{RG}B(G,R)<\infty$, and the assertion is true by Corollary 5.3. If $G$ is an $\mathrm{LH}\mathfrak{F}$-group of type $FP_{\infty}$, then by [23, Theorem A.2] and Corollary 4.2 we have $\mathrm{Gcd}_{R}G\leq\mathrm{Gcd}_{\mathbb{Z}}G<\infty$. ∎

### v5

@@ 1. Introduction
The origin of Gorenstein projective dimension may date back to the study of $G$-dimension by Auslander and Bridger [2] in 1960s. Let $G$ be any group. Over the group ring $RG$, the coefficient ring $R$ is an $RG$-module with trivial group action. The cohomological dimension ${\rm cd}_{R}G$, and the Gorenstein cohomological dimension ${\rm Gcd}_{R}G$, are defined as the projective dimension and Gorenstein projective dimension of the trivial $RG$-module $R$ respectively. Studying groups through their cohomological dimensions arose from both topological and algebraic sources; see [9]. We refer to Introduction of [22] for more details on the geometric significance of the Gorenstein cohomological dimension of groups.
@@
The “Gcd” can be considered as an assignment of invariants for the pairs of groups and coefficient rings $(G,R)$. In Section 3 and 4, we study the properties of Gcd under changes of groups and coefficient rings, respectively. we define an order $(H,S)\leq(G,R)$ for such pairs; see Definition 3.1. Inspired by [22, 39], we use Theorem 2.4 to show that when $R$ is a commutative Gorenstein regular ring, $\mathrm{Gcd}_{R}H\leq\mathrm{Gcd}_{R}G$ if $(H,R)\leq(G,R)$, and $\mathrm{Gcd}_{S}G\leq\mathrm{Gcd}_{R}G$ if $(G,S)\leq(G,R)$; see Proposition 3.2 and Proposition 4.1. These generalize [22, Propositions 2.1 and 2.4] and [39, Theorem 3.2]. Consequently, “Gcd” preserves the order of pairs of groups and commutative Gorenstein regular rings; see Corollary 4.3.
By Serre’s Theorem, ${\rm cd}_{\mathbb{Z}}G={\rm cd}_{\mathbb{Z}}H$ for a torsion-free group $G$ and any subgroup $H$ of finite index; see e.g. [9, Theorem VIII 3.1]. We have a Gorenstein version of Serre’s Theorem for any group and its subgroups of finite index over any coefficient ring; see Theorem 3.3. Note that the equality ${\rm Gcd}_{R}G={\rm Gcd}_{R}H$ was stated in [22, Corollary 2.10] under assumptions that the coefficient ring $R$ is of finite weak global dimension and $H$ is a normal subgroup of $G$.
@@
For model category $\mathcal{F}ib$, the associated homotopy category $\mathrm{Ho}(\mathcal{F}ib)$ is obtained by formally inverting the weak equivalences. It is standard that $\mathrm{Ho}(\mathcal{F}ib)$ is triangle equivalent to the stable category $\underline{\mathcal{C}of}(RG)$; see Corollary 6.4. Moreover, ${\rm Ho}(\mathcal{F}ib)$ is equivalent to the stable module category ${\rm StMod}(RG)$; see Theorem 6.8. It is worth to remark that the modules of ${\rm StMod}(RG)$ were assumed to be countably presented in [6], and Benson suggested that one can lift the countability hypothesis based on [6, Open Problem 2]. In fact, the objects of ${\rm StMod}(RG)$ are precisely the fibrant $RG$-modules, and the countability assumption of modules in ${\rm StMod}(RG)$ is not necessary.
In [33, Theorem 3.10], some equivalences of triangulated categories
@@ 2. Finiteness of Gorenstein cohomological dimension of groups
Let $M$ be a Gorenstein projective $RG$-module. Then there is a totally acyclic complex of projective $RG$-modules $\cdots\rightarrow P_{1}\rightarrow P_{0}\rightarrow P_{-1}\rightarrow\cdots$ such that $M\cong\mathrm{Ker}(P_{0}\rightarrow P_{-1})$. Since any projective $RG$-module is also $R$-projective, by restricting this totally acyclic complex, we get an acyclic complex of projective $R$-modules. Noting that $R$ is left Gorenstein regular, every projective $R$-module $P$ has finite injective dimension. Apply ${\rm Hom}_{R}(-,P)$ to the above complex, and it is easy to see that the acyclic complex of projective $R$-modules is totally acyclic by induction on the injective dimension of $P$. Hence, $M$ is also a Gorenstein projective $R$-module. ∎
@@
Let ${\rm Gcd}_{R}G={\rm Gpd}_{RG}R=n$. It follows from [29, Theorem 2.10] that there exists an exact sequence $0\rightarrow K\rightarrow M\rightarrow R\rightarrow 0$, where $M$ is a Gorenstein projective $RG$-module, and ${\rm pd}_{RG}K=n-1$. For $M$, there is an exact sequence of $RG$-modules $0\rightarrow M\rightarrow P\rightarrow L\rightarrow 0$, where $L$ is Gorenstein projective and $P$ is projective. We consider the following pushout of $M\rightarrow R$ and $M\rightarrow P$:
@@
From the middle row we infer that ${\rm pd}_{RG}\Lambda={\rm pd}_{RG}K+1=n$. It follows from Lemma 2.1 that $L$ is also a Gorenstein projective $R$-module, and then the sequence $0\rightarrow R\rightarrow\Lambda\rightarrow L\rightarrow 0$ is $R$-split. Moreover, as an $R$-module, $\Lambda\cong L\oplus R$ is Gorenstein projective. By [26, Proposition 10.2.3], which says that projective dimension of any Gorenstein projective module is either zero or infinity, we induce from ${\rm pd}_{R}\Lambda\leq{\rm pd}_{RG}\Lambda=n$ that $\Lambda$ is a projective $R$-module. This completes the proof. ∎
@@
It is well known that for any finite group $G$ and any coefficient ring $R$, ${\rm Gcd}_{R}G=0$; see [1, Proposition 2.19] for $R=\mathbb{Z}$, [22, Corollary 2.3] for any commutative ring $R$, and Proposition 3.4 below for any associative ring $R$.
@@ 3. Gorenstein cohomological dimension under changes of groups
By Serre’s theorem, if $G$ is a torsion-free group and $H$ is a subgroup of finite index, then $\mathrm{cd}_{\mathbb{Z}}H=\mathrm{cd}_{\mathbb{Z}}G$; see e.g. [37, Theorem 9.2] or [9, Theorem VIII 3.1]. The following result might be regarded as a Gorenstein version of Serre’s theorem. We remark that the equality for the coefficient ring $\mathbb{Z}$ holds by [1, Proposition 2.18], and the result was also proved in [22, Corollary 2.10] under the assumptions that the coefficient ring $R$ is of finite weak global dimension and $H$ is a normal subgroup of $G$.
@@
that $\mathrm{Ext}_{RH}^{1}(L_{0},Q)=0$. Consider $0\rightarrow K\stackrel{{\scriptstyle d}}{{\rightarrow}}P_{0}\rightarrow L_{0}\rightarrow 0$ as an exact sequence of $RH$-modules, where $K$ is Gorenstein projective and $P_{0}$ is projective. By [29, Corollary 2.11], $L_{0}$ is a Gorenstein projective $RH$-module. Hence, for any projective $RG$-module $P$, we infer from ${\rm Ext}^{1}_{RG}(L_{0},{\rm Ind}_{H}^{G}P)\cong{\rm Ext}_{RH}^{1}(L_{0},P)=0$ that ${\rm Ext}^{1}_{RG}(L_{0},P)=0$. This implies that the sequence $0\rightarrow K\rightarrow P_{0}\rightarrow L_{0}\rightarrow 0$ remains exact after applying ${\rm Hom}_{RG}(-,P)$.
Then, repeat the above argument for $L_{0}$, we will obtain inductively an acyclic complex $0\rightarrow K\rightarrow P_{0}\rightarrow P_{-1}\rightarrow P_{-2}\rightarrow\cdots$ with each $P_{i}$ a projective $RG$-module, which remains acyclic after applying ${\rm Hom}_{RG}(-,P)$ for any projective $RG$-module $P$. Recall that ${\rm Ext}^{i}_{RG}(K,P)=0$. Consequently, we infer from [29, Proposition 2.3] that $K$ is a Gorenstein projective $RG$-module, as required. ∎
@@
(2$\Longrightarrow$(3) is trivial, and (1$\Longleftrightarrow$(3) follows from [22, Corollary 2.3].
@@
Let $G$ be a group, $H$ be a finite subgroup of $G$, and denote by $N_{G}(H)$ the normalizer of $H$ in $G$. The Weyl group of $H$ is $W=N_{G}(H)/H$. Note that any module over the group $G$ on which the subgroup $H$ acts trivially induces a module over the Weyl group $W$.
@@
Let $R$ be a commutative Gorenstein regular ring, $G$ be a group. For any finite subgroup $H$ of $G$, we have ${\rm Gcd}_{R}(N_{G}(H)/H)\leq{\rm Gcd}_{R}G$.
@@ 4. Extension of coefficient rings
Since the inequality obviously holds if $\mathrm{Gcd}_{R}G=\infty$, it only suffices to consider the case where $\mathrm{Gcd}_{R}G=n$ is finite. In this case, it follows immediately from Lemma 2.2 that there exists an $R$-split monomorphism of $RG$-modules $\iota:R\rightarrow\Lambda$, where $\Lambda$ is $R$-projective and $\mathrm{pd}_{RG}\Lambda=n$. Note that $S$ is an extension ring of $R$, then $S$ is both a trivial $SG$-module and an $R$-module. By applying $S\otimes_{R}-$, we have an $S$-split monomorphism of $SG$-modules $S\otimes_{R}\iota:S\rightarrow S\otimes_{R}\Lambda$, where $S\otimes_{R}\Lambda$ is $S$-projective since $\Lambda$ is $R$-projective.
@@
Since $S\otimes_{R}P_{i}$ are projective $SG$-modules, we have $\mathrm{pd}_{SG}(S\otimes_{R}\Lambda)\leq n$. For the trivial $SG$-module $S$, it follows immediately from [22, Proposition 1.4] that $\mathrm{Gpd}_{SG}S\leq\mathrm{pd}_{SG}(S\otimes_{R}\Lambda)$. Then, $\mathrm{Gcd}_{S}G\leq\mathrm{Gcd}_{R}G$ holds as required. ∎
Specifically, considering the pairs $(G,R)\leq(G,\mathbb{Z})$ and $(G,\mathbb{Q})\leq(G,\mathbb{Z})$, we reobtain [22, Propostion 2.1] and [39, Theorem 3.2] respectively.
@@
If $(H,S)\leq(G,R)$, then we have both $(H,S)\leq(G,S)$ and $(G,S)\leq(G,R)$. By Proposition 3.2, it follows that ${\rm Gcd}_{S}H\leq{\rm Gcd}_{S}G$. By Proposition 4.1, $\mathrm{Gcd}_{S}G\leq\mathrm{Gcd}_{R}G$ holds. Then, we deduce the required inequality $\mathrm{Gcd}_{S}H\leq\mathrm{Gcd}_{R}G$. ∎
@@
Note that the second inequality follows from Corollary 4.2. The first inequality obviously holds if ${\rm Gcd}_{R}G$ is infinite. Furthermore, if we assume the finiteness of ${\rm Gcd}_{R}G$, then the inequality follows immediately from Lemmas 2.2 and 2.3. ∎
@@ 5. Cofibrant modules and Gorenstein projective modules
Recall that a group $G$ is of type $\Phi_{R}$, if it has the property that an $RG$-module is of finite projective dimension if and only if its restriction to any finite subgroup is of finite projective dimension; see [33, Definition 2.1]. It was conjectured that a group $G$ is of type $\Phi_{\mathbb{Z}}$ if and only if $G$ admits a finite dimensional model for $\underline{E}G$, the classifying space for the family of the finite subgroups of $G$; the sufficient condition was proved to be true. We refer to [38] for the details.
@@
Let $G$ be an $\mathrm{LH}\mathfrak{F}$-group, $R$ a commutative Gorenstein regular ring. For any Gorenstein projective $RG$-module $M$, if ${\rm pd}_{R}M<\infty$, then $M$ is cofibrant.
@@
We might get the following, which recovers [19, Corollary C], [7, Theorem 3.11] and [8, Theorem 3.4].
@@
Proof.
Invoking Proposition 5.2, it suffices to prove that for any $RG$-module $M$, $\mathrm{pd}_{RG}M\otimes_{R}B(G,R)\leq\mathrm{Gpd}_{RG}M$. Since the inequality which is to be proved is obvious if $\mathrm{Gpd}_{RG}M=\infty$, it only suffices to consider the case where $\mathrm{Gpd}_{RG}M=n$ is finite. Then, there exists an exact sequence of $RG$-modules
$0\longrightarrow M_{n}\longrightarrow\cdots\longrightarrow M_{1}\longrightarrow M_{0}\longrightarrow M\longrightarrow 0,$
where $M_{0},M_{1},\cdots,M_{n}$ are Gorenstein projective; they are also cofibrant modules by Propositions 5.6 and 5.7. By applying $-\otimes_{R}B(G,R)$ to this sequence, we infer that $\mathrm{pd}_{RG}M\otimes_{R}B(G,R)\leq n$ since $M_{i}\otimes_{R}B(G,R)$ are projective $RG$-modules. ∎
@@
If $G$ is of type $\Phi_{R}$, then by [6, Lemma 3.4(ii)] we have $\mathrm{pd}_{RG}B(G,R)<\infty$, and the assertion is true by Corollary 5.3. If $G$ is an $\mathrm{LH}\mathfrak{F}$-group of type $FP_{\infty}$, then by [23, Theorem A.2] we have $\mathrm{Gcd}_{\mathbb{Z}}G<\infty$. Moreover, by Corollaries 4.2 and 5.8 we have $\mathrm{pd}_{RG}B(G,R)=\mathrm{Gcd}_{R}G\leq\mathrm{Gcd}_{\mathbb{Z}}G<\infty$. ∎